# Practical solutions to the relative pose of three calibrated cameras

### v4

@@ 1 Introduction
00footnotetext: ∗ Equal contributionCamera geometry estimation is crucial in many computer vision applications, e.g., visual navigation [49], Structure-from-Motion [53], augmented reality [4], self-driving cars [17], and visual localization [48]. Due to noise and outliers in input correspondences, the predominant way for camera geometry estimation is to use a hypothesis-and-test framework, e.g., RANSAC [15, 7, 46, 2]. For RANSAC-like methods, using as few (ideally the minimal number of) correspondences as possible for estimation is important since the number of RANSAC iterations (and thus its run-time) grows exponentially with the number of correspondences required for the model estimation.
@@
Minimal camera geometry problems often result in complex systems of polynomial equations. Efficient algebraic methods helped solve many previously unsolved problems [55, 3, 30, 26, 27, 56]. Still, they fail to generate efficient and/or numerically stable solutions for some configurations. In this paper, we study one such challenging problem: Estimating the relative pose of three calibrated cameras. This problem has received attention for a long time [20, 45, 32, 35, 1]. However, due to its complexity, it is still not considered fully solved. There are no efficient and practical solutions for most of the minimal configurations of point and/or line correspondences [25]. One such configuration that is particularly interesting is the notoriously difficult configuration of four points in three calibrated views [45, 36], known as the 4p3v problem.
@@
State-of-the-art algebraic and numerical methods are known to fail in generating efficient and numerically stable solutions to the 4p3v problem. The existing methods for solving this problem are only approximate [22, 37]. By solving only for one (or a few) solutions from the 272 solutions of the 4p3v problem [22], and by discretely sampling the space of potential solutions [37], the existing 4p3v methods can often fail, i.e., the returned solution can be, in general, arbitrarily far from the geometrically correct solution. To decrease the failure rate, both methods [22, 37] require a lot of tuning and are challenging to re-implement.111There is no publicly available implementation for [37]. Implementation of [22] is quite complex and requires a non-negligible effort to run.
@@
In contrast to complex solutions to the configuration of four points in three views, there is a simple and efficient solution for the configuration where five correspondences are detected in the first two views and three of these points are also visible in the third view. In this case, the 5pt+P3P solver first estimates the relative pose of two cameras from five correspondences using an efficient 5pt solver [36], and then registers the third camera using a P3P solver [41].
@@
Following the idea of first estimating the relative pose of two cameras, we propose two novel approaches for solving the 4p3v problem. Our solutions are based on the simple idea of using four correspondences to estimate an approximate geometry of the first two views and then registering the third camera using a P3P solver [41].
@@
In the first approach, we approximate the geometry of the first two views using affine cameras. The approximation using affine cameras was shown to provide good accuracy for relative pose estimation if coupled with geometry refitting and local optimization inside RANSAC [42]. We denote the proposed affine-based 4p3v solver as 4p3v(A).
@@
In the second approach, we use a less restrictive approximation. In this case, we estimate the full perspective geometry, i.e., the full 5DoF essential matrix, using four input correspondences and one additional approximate correspondence in two views. We generate the approximate correspondence using the locations of three of the four input correspondences. Under the assumption of a para-perspective projection, i.e., of affine geometry, the mean point of three 3D points is projected to the mean points in both images [59]. Thus, the new correspondence is generated as the correspondence between the mean points of three corresponding points detected in two views. This approximate mean-point correspondence can also be seen as a correspondence under the -order approximation of the homography defined by the plane passing through the three corresponding 3D points. We denote the proposed mean point-based 4p3v solver as 4p3v(M).
@@
While the proposed 4p3v(M) solver returns significantly more accurate poses than the 4p3v(A) solver, both solvers may not be accurate enough if not properly treated inside RANSAC. Thus, in this paper, we propose several ways to improve the accuracy of the proposed approximate solvers. (i) To compensate for noise in the mean point correspondences, we introduce the 4p3v(M) solver that generates two additional correspondences in the vicinity of the mean point, and in the first step calls the 5pt solver [36] three times. (ii) To improve the approximate geometry estimated between the first two views, we refit this geometry using a non-minimal relative pose solver and inliers obtained from the approximate geometry (ENM). (iii) We use the fourth correspondence in the third view to filter out geometrically infeasible solutions (+F) and to refine the solutions on four input correspondences in three views using just a few iterations of Levenberg-Marquardt (LM) refinement (+R). While conceptually simple and efficient, the novel solvers achieve state-of-the-art results on real data.
@@
We propose two groups of novel solutions for the well-known and challenging 4p3v problem: 4p3v(A)- and 4p3v(M)-based solvers. These solutions are based on the simple idea of using four correspondences to estimate an approximate geometry in the first two views. Compared to state-of-the-art 4p3v solvers [22, 37], which are non-trivial and difficult to re-implement such that they are numerically stable and fast, our new solutions can be easily implemented using existing efficient implementations of the linear 4pt affine fundamental matrix solver [18], the 5pt solver [36] and the P3P solver [41]. The source code is available at https://github.com/kocurvik/threeview.
@@ 2 Related work
Estimating the relative pose of three cameras from a minimal number of point and line correspondences is known as an extremely challenging problem. For three uncalibrated cameras, 6 point correspondences are necessary to estimate the trifocal tensor, with a solution known for a long time [44, 57]. Solutions to three minimal combinations of points and lines are presented in [39]. The configuration of 9 lines is more challenging and was solved only recently [29]. Yet, the solver is far from practical and runs 17.8s.
@@
For calibrated cameras, the configuration that attracts most of the attention is the configuration of four points in three views (the 4p3v problem). Note that this is not a minimal configuration since it generates 12 constraints for 11 degrees-of-freedom (DoF). The 4p3v problem is known to be extremely difficult to solve. Several papers present mostly theoretical results [32, 35, 1]. For four triplets of exact points without noise, it is shown that the 4p3v problem has, in general, a unique solution [20, 45].
@@
To the best of our knowledge, there are only two reasonably efficient solutions to the 4p3v problem reported in the literature. The first solver [37] is based on a similar idea as our solvers, i.e., to first estimate the relative pose of two cameras and then register the third camera using a P3P solver [41]. To compute the pose of the first two cameras using information only from four point correspondences, the solver needs additional information about the position of one epipole. The paper shows that the four point correspondences between two calibrated views constrain the epipole in each image to lie on a curve of degree ten. Thus, the solver performs one-dimensional exhaustive search and sweeps a -degree curve of possible epipoles. For each potential epipole, it computes the relative pose of two cameras, registers the third camera using three triangulated points, and finally extracts the solution minimizing the reprojection error of the fourth point in the third view. Evaluation of the solver on one potential epipole is fast. Yet, in contrast to our proposed 4p3v(M)-based solvers, the error of the sampled epipole for one fixed point on the curve is not bounded since the true epipole can lie anywhere on the curve, and in many scenarios it is very far from the image center (e.g., outside the image). Thus, to obtain reasonably accurate and stable results, usually 1,000 candidates need to be evaluated. Even then, refinement at multiple local minima is required to improve the accuracy. The runtimes reported for this solver were depending on the number of points searched. On a small number of synthetic experiments, the paper shows that the translation error returned by the proposed solver is usually higher than the error returned by the 5pt+P3P solver. The literature does not compare against [37] as there is no publicly available implementation and it is hard to re-implement.
@@
The second efficient solver to the 4p3v problem was published only recently [22]. In this paper, the authors first transform the 4p3v problem into a minimal problem by considering a line passing through the last correspondence in the third view. The resulting system of equations is solved using an efficient Homotopy continuation (HC) method [14, 54]. To avoid computing large numbers of spurious solutions, an MLP-based classifier is trained. For a given problem , it selects one or several starting problem-solution pairs (so-called anchors), such that the geometrically meaningful/correct solution of can be obtained by HC starting from this anchor. This strategy is fast, running on average per solution. However, it has a high failure rate. The success rate of the 4p3v solver reported in [22] on two test datasets and data without noise is . [22] do not show results for a real scenario, i.e., a RANSAC-like framework with noisy data. Providing such an evaluation, we show that our much simpler solvers outperform [22].
@@
Solutions to the 4p3v problem for orthographic and para-perspective views were presented in [58, 34]. In [34], the author suggested an iterative approach for updating to perspective views, but reported results only on a few synthetic instances. According to our extensive experiments, the update does not work on real data with general perspective cameras. This solver is returning large errors even after incorporating it into RANSAC with local optimization. We see two main reasons: (i) The set of inliers that satisfy an approximate para-perspective camera model in all three images is usually quite small, and (ii) the estimated model is often very far from the perspective optimum, and thus local optimization does not converge to a good solution.
@@
In [13, 25, 12], the authors aim to classify and derive the number of solutions for different minimal configurations of points and lines in three calibrated views. Solutions to two minimal configurations combining points and lines were proposed in [14] and solved using a HC method [54]. Due to their complexity, the solvers are not practical. A GPU HC method was also used to solve minimal problems of four points/six lines in three views for a generalized camera in [11] and of four points in three cameras with an unknown shared focal length (4p3vf) in [5, 8]. Efficient GPU implementations of the 4p3vf solvers run to . These times are still too slow for practical applications.
@@
An approximation of perspective cameras using affine ones was used in several papers on camera geometry estimation [9]. [38] and [21] propose iterative approaches to the absolute camera pose estimation problem, i.e., the PnP problem. Both methods first compute the pose for an affine camera (weak- or para-perspective). Then the error induced by the affine camera approximation is used to adjust the constraints on the pose and recompute it. These methods do not guarantee convergence to the perspective model solution. In [42], affine cameras are used to efficiently solve the relative pose problem of two uncalibrated cameras from two affine correspondences. These affine correspondences are transformed to six point or two ellipse correspondences and then used to estimate the affine fundamental matrix. The paper also proposes a RANSAC framework that uses local optimization [31] to estimate the full fundamental matrix using inliers from the approximate affine model.
@@
Points that are sampled based on feature geometry to generate point correspondences from affine or scale- and orientation-invariant feature correspondences were also used in [40, 16, 43]. In contrast, our M-based solvers use only point correspondences, without associated feature geometry, to generate an additional point correspondence.
@@ 3 Estimating the relative pose of three cameras
In this section, we describe different solutions for estimating the relative pose of three calibrated cameras. We start with a baseline solution for the minimal configuration of three points visible in all three cameras and two additional points visible in two of the three cameras (the (5,5,3) configuration). Next, we present our novel solutions for the configuration of four points visible in all three cameras (the (4,4,4) configuration). This configuration generates an over-constrained problem. In this case, we have one more constraint than DoF. A minimal solution would need to drop one constraint, e.g., by considering only a line passing through one of the points in the third view [22] or by considering a “half" point correspondence. Since, in practice, we always have full correspondences and sampling one less point in one view leads to an under-constrained problem, the (4,4,4) configuration, is usually considered “minimal".
5pt+P3P solver: The 5pt+P3P solver first estimates the relative pose of two cameras from 5 image point correspondences using the efficient 5pt solver [36]. Next, the three points visible in all three views are triangulated. Finally, the third camera is registered using the three 2D-3D point correspondences and the well-known efficient P3P solver [41].
This straightforward solver, which is based on existing efficient solvers [36, 41], was discussed in several works [13, 36, 47, 37]. [37] showed that the 5pt+P3P solver performs better than their dedicated 4p3v solver on synthetic data. Yet, the most recent works [14, 22] that study the three view relative pose problem do not discuss the 5pt+P3P solver and do not use it as a baseline for comparison. To the best of our knowledge, the performance of this solver on real data and within state-of-the-art RANSAC frameworks in the context of the 4p3v problem has not been extensively studied. Our paper fills this gap in the literature.

### v5

@@ 1 Introduction
00footnotetext: ∗ Equal contributionCamera geometry estimation is crucial in many computer vision applications, e.g., visual navigation [DBLP:journals/ram/ScaramuzzaF11], Structure-from-Motion [Snavely-IJCV-2008], augmented reality [Castle08ISWC], self-driving cars [hane20173d], and visual localization [Sattler16PAMI]. Due to noise and outliers in input correspondences, the predominant way for camera geometry estimation is to use a hypothesis-and-test framework, e.g., RANSAC [Fischler-Bolles-ACM-1981, Chum-2003, DBLP:journals/pami/RaguramCPMF13, barath2017graph]. For RANSAC-like methods, using as few (ideally the minimal number of) correspondences as possible for estimation is important since the number of RANSAC iterations (and thus its run-time) grows exponentially with the number of correspondences required for the model estimation.
@@
Minimal camera geometry problems often result in complex systems of polynomial equations. Efficient algebraic methods helped solve many previously unsolved problems [Stewenius-CVPR-2005, bujnak_cvpr2008, larsson2019revisiting, DBLP:conf/cvpr/KukelovaP07, kukelova2013real, Stewenius-ISPRS-2006]. Still, they fail to generate efficient and/or numerically stable solutions for some configurations. In this paper, we study one such challenging problem: Estimating the relative pose of three calibrated cameras. This problem has received attention for a long time [holt1995, Quan2006, leonardos_cvpr2015, Martyushev16, Aholt2014]. However, due to its complexity, it is still not considered fully solved. There are no efficient and practical solutions for most of the minimal configurations of point and/or line correspondences [Kileel2017]. One such configuration that is particularly interesting is the notoriously difficult configuration of four points in three calibrated views [Quan2006, Nister-5pt-PAMI-2004], known as the 4p3v problem.
@@
State-of-the-art algebraic and numerical methods are known to fail in generating efficient and numerically stable solutions to the 4p3v problem. The existing methods for solving this problem are only approximate [Hruby_cvpr2022, DBLP:journals/ijcv/NisterS06]. By solving only for one (or a few) solutions from the 272 solutions of the 4p3v problem [Hruby_cvpr2022], and by discretely sampling the space of potential solutions [DBLP:journals/ijcv/NisterS06], the existing 4p3v methods can often fail, i.e., the returned solution can be, in general, arbitrarily far from the geometrically correct solution. To decrease the failure rate, both methods [Hruby_cvpr2022, DBLP:journals/ijcv/NisterS06] require a lot of tuning and are challenging to re-implement.111There is no publicly available implementation for [DBLP:journals/ijcv/NisterS06]. Implementation of [Hruby_cvpr2022] is quite complex and requires a non-negligible effort to run.
@@
In contrast to complex solutions to the configuration of four points in three views, there is a simple and efficient solution for the configuration where five correspondences are detected in the first two views and three of these points are also visible in the third view. In this case, the 5pt+P3P solver first estimates the relative pose of two cameras from five correspondences using an efficient 5pt solver [Nister-5pt-PAMI-2004], and then registers the third camera using a P3P solver [lambda-twist].
@@
Following the idea of first estimating the relative pose of two cameras, we propose two novel approaches for solving the 4p3v problem. Our solutions are based on the simple idea of using four correspondences to estimate an approximate geometry of the first two views and then registering the third camera using a P3P solver [lambda-twist].
@@
In the first approach, we approximate the geometry of the first two views using affine cameras. The approximation using affine cameras was shown to provide good accuracy for relative pose estimation if coupled with geometry refitting and local optimization inside RANSAC [pritts_ivcnz13]. We denote the proposed affine-based 4p3v solver as 4p3v(A).
@@
In the second approach, we use a less restrictive approximation. In this case, we estimate the full perspective geometry, i.e., the full 5DoF essential matrix, using four input correspondences and one additional approximate correspondence in two views. We generate the approximate correspondence using the locations of three of the four input correspondences. Under the assumption of a para-perspective projection, i.e., of affine geometry, the mean point of three 3D points is projected to the mean points in both images [Zhang2014]. Thus, the new correspondence is generated as the correspondence between the mean points of three corresponding points detected in two views. This approximate mean-point correspondence can also be seen as a correspondence under the -order approximation of the homography defined by the plane passing through the three corresponding 3D points. We denote the proposed mean point-based 4p3v solver as 4p3v(M).
@@
While the proposed 4p3v(M) solver returns significantly more accurate poses than the 4p3v(A) solver, both solvers may not be accurate enough if not properly treated inside RANSAC. Thus, in this paper, we propose several ways to improve the accuracy of the proposed approximate solvers. (i) To compensate for noise in the mean point correspondences, we introduce the 4p3v(M) solver that generates two additional correspondences in the vicinity of the mean point, and in the first step calls the 5pt solver [Nister-5pt-PAMI-2004] three times. (ii) To improve the approximate geometry estimated between the first two views, we refit this geometry using a non-minimal relative pose solver and inliers obtained from the approximate geometry (ENM). (iii) We use the fourth correspondence in the third view to filter out geometrically infeasible solutions (+F) and to refine the solutions on four input correspondences in three views using just a few iterations of Levenberg-Marquardt (LM) refinement (+R). While conceptually simple and efficient, the novel solvers achieve state-of-the-art results on real data.
@@
We propose two groups of novel solutions for the well-known and challenging 4p3v problem: 4p3v(A)- and 4p3v(M)-based solvers. These solutions are based on the simple idea of using four correspondences to estimate an approximate geometry in the first two views. Compared to state-of-the-art 4p3v solvers [Hruby_cvpr2022, DBLP:journals/ijcv/NisterS06], which are non-trivial and difficult to re-implement such that they are numerically stable and fast, our new solutions can be easily implemented using existing efficient implementations of the linear 4pt affine fundamental matrix solver [hartley2006multiple], the 5pt solver [Nister-5pt-PAMI-2004] and the P3P solver [lambda-twist]. The source code is available at https://github.com/kocurvik/threeview and https://doi.org/10.5281/zenodo.16599943. The data used are available and can be accessed at https://doi.org/10.5281/zenodo.16603086.
@@ 2 Related work
Estimating the relative pose of three cameras from a minimal number of point and line correspondences is known as an extremely challenging problem. For three uncalibrated cameras, 6 point correspondences are necessary to estimate the trifocal tensor, with a solution known for a long time [Quan_pami95, Torr97a]. Solutions to three minimal combinations of points and lines are presented in [Oskarsson_bmvc2004]. The configuration of 9 lines is more challenging and was solved only recently [larsson2017efficient]. Yet, the solver is far from practical and runs 17.8s.
@@
For calibrated cameras, the configuration that attracts most of the attention is the configuration of four points in three views (the 4p3v problem). Note that this is not a minimal configuration since it generates 12 constraints for 11 degrees-of-freedom (DoF). The 4p3v problem is known to be extremely difficult to solve. Several papers present mostly theoretical results [leonardos_cvpr2015, Martyushev16, Aholt2014]. For four triplets of exact points without noise, it is shown that the 4p3v problem has, in general, a unique solution [holt1995, Quan2006].
@@
To the best of our knowledge, there are only two reasonably efficient solutions to the 4p3v problem reported in the literature. The first solver [DBLP:journals/ijcv/NisterS06] is based on a similar idea as our solvers, i.e., to first estimate the relative pose of two cameras and then register the third camera using a P3P solver [lambda-twist]. To compute the pose of the first two cameras using information only from four point correspondences, the solver needs additional information about the position of one epipole. The paper shows that the four point correspondences between two calibrated views constrain the epipole in each image to lie on a curve of degree ten. Thus, the solver performs one-dimensional exhaustive search and sweeps a -degree curve of possible epipoles. For each potential epipole, it computes the relative pose of two cameras, registers the third camera using three triangulated points, and finally extracts the solution minimizing the reprojection error of the fourth point in the third view. Evaluation of the solver on one potential epipole is fast. Yet, in contrast to our proposed 4p3v(M)-based solvers, the error of the sampled epipole for one fixed point on the curve is not bounded since the true epipole can lie anywhere on the curve, and in many scenarios it is very far from the image center (e.g., outside the image). Thus, to obtain reasonably accurate and stable results, usually 1,000 candidates need to be evaluated. Even then, refinement at multiple local minima is required to improve the accuracy. The runtimes reported for this solver were depending on the number of points searched. On a small number of synthetic experiments, the paper shows that the translation error returned by the proposed solver is usually higher than the error returned by the 5pt+P3P solver. The literature does not compare against [DBLP:journals/ijcv/NisterS06] as there is no publicly available implementation and it is hard to re-implement.
@@
The second efficient solver to the 4p3v problem was published only recently [Hruby_cvpr2022]. In this paper, the authors first transform the 4p3v problem into a minimal problem by considering a line passing through the last correspondence in the third view. The resulting system of equations is solved using an efficient Homotopy continuation (HC) method [Fabbri_CVPR2020, SommeseAndrewJ2005Tnso]. To avoid computing large numbers of spurious solutions, an MLP-based classifier is trained. For a given problem , it selects one or several starting problem-solution pairs (so-called anchors), such that the geometrically meaningful/correct solution of can be obtained by HC starting from this anchor. This strategy is fast, running on average per solution. However, it has a high failure rate. The success rate of the 4p3v solver reported in [Hruby_cvpr2022] on two test datasets and data without noise is . [Hruby_cvpr2022] do not show results for a real scenario, i.e., a RANSAC-like framework with noisy data. Providing such an evaluation, we show that our much simpler solvers outperform [Hruby_cvpr2022].
@@
Solutions to the 4p3v problem for orthographic and para-perspective views were presented in [xu_ortho4p3v, Higgins-4p3v91]. In [Higgins-4p3v91], the author suggested an iterative approach for updating to perspective views, but reported results only on a few synthetic instances. According to our extensive experiments, the update does not work on real data with general perspective cameras. This solver is returning large errors even after incorporating it into RANSAC with local optimization. We see two main reasons: (i) The set of inliers that satisfy an approximate para-perspective camera model in all three images is usually quite small, and (ii) the estimated model is often very far from the perspective optimum, and thus local optimization does not converge to a good solution.
@@
In [Duff_PL1P, Kileel2017, duff2019plmp], the authors aim to classify and derive the number of solutions for different minimal configurations of points and lines in three calibrated views. Solutions to two minimal configurations combining points and lines were proposed in [Fabbri_CVPR2020] and solved using a HC method [SommeseAndrewJ2005Tnso]. Due to their complexity, the solvers are not practical. A GPU HC method was also used to solve minimal problems of four points/six lines in three views for a generalized camera in [Ding_2023_ICCV] and of four points in three cameras with an unknown shared focal length (4p3vf) in [Chien_2022_CVPR, Cin_2024_CVPR]. Efficient GPU implementations of the 4p3vf solvers run to . These times are still too slow for practical applications.
@@
An approximation of perspective cameras using affine ones was used in several papers on camera geometry estimation [Collins2014InfinitesimalPP]. [oberkampf1996] and [Horaud97] propose iterative approaches to the absolute camera pose estimation problem, i.e., the PnP problem. Both methods first compute the pose for an affine camera (weak- or para-perspective). Then the error induced by the affine camera approximation is used to adjust the constraints on the pose and recompute it. These methods do not guarantee convergence to the perspective model solution. In [pritts_ivcnz13], affine cameras are used to efficiently solve the relative pose problem of two uncalibrated cameras from two affine correspondences. These affine correspondences are transformed to six point or two ellipse correspondences and then used to estimate the affine fundamental matrix. The paper also proposes a RANSAC framework that uses local optimization [lebedaLO] to estimate the full fundamental matrix using inliers from the approximate affine model.
@@
Points that are sampled based on feature geometry to generate point correspondences from affine or scale- and orientation-invariant feature correspondences were also used in [perd2006epipolar, BEEP, pritts-cvpr2018]. In contrast, our M-based solvers use only point correspondences, without associated feature geometry, to generate an additional point correspondence.
@@ 3 Estimating the relative pose of three cameras
In this section, we describe different solutions for estimating the relative pose of three calibrated cameras. We start with a baseline solution for the minimal configuration of three points visible in all three cameras and two additional points visible in two of the three cameras (the (5,5,3) configuration). Next, we present our novel solutions for the configuration of four points visible in all three cameras (the (4,4,4) configuration). This configuration generates an over-constrained problem. In this case, we have one more constraint than DoF. A minimal solution would need to drop one constraint, e.g., by considering only a line passing through one of the points in the third view [Hruby_cvpr2022] or by considering a “half” point correspondence. Since, in practice, we always have full correspondences and sampling one less point in one view leads to an under-constrained problem, the (4,4,4) configuration, is usually considered “minimal”.
5pt+P3P solver: The 5pt+P3P solver first estimates the relative pose of two cameras from 5 image point correspondences using the efficient 5pt solver [Nister-5pt-PAMI-2004]. Next, the three points visible in all three views are triangulated. Finally, the third camera is registered using the three 2D-3D point correspondences and the well-known efficient P3P solver [lambda-twist].
This straightforward solver, which is based on existing efficient solvers [Nister-5pt-PAMI-2004, lambda-twist], was discussed in several works [Duff_PL1P, Nister-5pt-PAMI-2004, Rodehorst2017, DBLP:journals/ijcv/NisterS06]. [DBLP:journals/ijcv/NisterS06] showed that the 5pt+P3P solver performs better than their dedicated 4p3v solver on synthetic data. Yet, the most recent works [Fabbri_CVPR2020, Hruby_cvpr2022] that study the three view relative pose problem do not discuss the 5pt+P3P solver and do not use it as a baseline for comparison. To the best of our knowledge, the performance of this solver on real data and within state-of-the-art RANSAC frameworks in the context of the 4p3v problem has not been extensively studied. Our paper fills this gap in the literature.